OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6. A: Y=1/3x - 8 B: Y=-3x+12 C: Y=3x-24 D: Y=-1/3x-4

OpenStudy (anonymous):

perpendicular means the slopes are negative reciprocals of each other. an easier way is to multiply the slopes together and set the product equal to -1. solve for the one you need. Then use the point-slope form amd the given point to find the needed equation.

OpenStudy (anonymous):

So multiply 3x by the slopes in the answers till i get one? Sorry I don't quite understand.

OpenStudy (anonymous):

line 1 has a slope of 3. call this\[m _{1}\] let \[m _{2}\]be the slope of the line you need to find the equation of. Then we have \[m _{1}\times m _{2}=-1 \implies m _{2}= \frac{ -1 }{ m _{1} } \implies m _{2}= \frac{ -1 }{ 3 }\]

OpenStudy (anonymous):

Oh, I see. But how did you get the -1? Does that stay the same when using this method or does it change?

OpenStudy (anonymous):

if two lines are perpendicular, the product of their slopes will always be -1.

OpenStudy (anonymous):

ok, thank you.